# Filamentary Ejecta Network in Cassiopeia A Reveals Fingerprints of the Supernova Explosion Mechanism

### v1

@@ Appendix B Effects of the magnetic field on the structure of unshocked ejecta
While the magnetic field has little influence on unshocked ejecta, it plays a crucial role in the dynamics of shocked ejecta. In particular, it suppresses the growth of HD instabilities (such as Rayleigh-Taylor and Kelvin-Helmholtz instabilities) at the contact discontinuity between shocked ejecta and the shocked CSM and can enhance the efficiency of radiative cooling (e.g., Orlando et al. 2008). In a companion paper (Orlando et al., submitted to A&A), we investigate the structure of a shocked, dense CSM shell in front of Cas A recently explored with JWST and termed the “Green Monster” (Milisavljevic et al. 2024; De Looze et al. 2024). In this separate work we examine in detail the role of the magnetic field in shaping newly discovered hole and ring features and, more generally, the structure of the mixing region between the forward and reverse shocks.
@@ Appendix C Online multi-media material
Interactive 3D Graphics. These models are accessible online through the sketchfab website, offering an interactive experience - even through virtual reality device (3DMAP-VR project; Orlando et al. 2019b, 2023) - to complement the figures in the paper:

### v2

@@ Appendix B Effects of the magnetic field on the structure of unshocked ejecta
While the magnetic field has little influence on unshocked ejecta, it plays a crucial role in the dynamics of shocked ejecta. In particular, it suppresses the growth of HD instabilities (such as Rayleigh-Taylor and Kelvin-Helmholtz instabilities) at the contact discontinuity between shocked ejecta and the shocked CSM and can enhance the efficiency of radiative cooling (e.g., Orlando et al. 2008). In a companion paper (Orlando et al. 2025), we investigate the structure of a shocked, dense CSM shell in front of Cas A recently explored with JWST and termed the “Green Monster” (Milisavljevic et al. 2024; De Looze et al. 2024). In this separate work we examine in detail the role of the magnetic field in shaping newly discovered hole and ring features and, more generally, the structure of the mixing region between the forward and reverse shocks.
@@ Appendix C Online multi-media material
Interactive 3D Graphics. These models are available as supplementary material on the A&A webpage and accessible online through the sketchfab website, offering an interactive experience - even through virtual reality device (3DMAP-VR project; Orlando et al. 2019b, 2023) - to complement the figures in the paper: